# A greedy heuristic for graph burning

### v1

@@ 6 Discussion
GBP is an NP-hard minimization problem with application to information spreading problems. Some heuristics and metaheuristics have been proposed to find feasible solutions. However, research about more straightforward greedy approaches wasn’t fully addressed. This paper introduces a simple deterministic greedy heuristic, Gr (Algorithm 3), tailored for the CMCP, a problem to which GBP is reduced if $b(G)$ is known in advance; the issue of not knowing this value is tackled by using binary search (Algorithm 4). Although Gr cannot guarantee to find optimal solutions for paths and cycles for which exact $\mathcal{O}(n)$ algorithms exist, it performed unexpectedly well over real-world and synthetic graphs. Namely, for a benchmark dataset of small real-world networks, Gr found 31 optimal solutions out of 40, each in no more than 0.234 seconds. Its plus version, GrP, found 38 out of 40 optimal solutions, each in less than 2.4 minutes (See Table 2). Regarding larger real-world networks, Gr found 7 out of 9 optimal solutions and nine new best-known solutions, each in less than 426 seconds ($\sim$7 minutes); GrP found 8 out of 9 optimal solutions, though it required up to 46798 seconds ($\sim$13 hours) (See Table 3). Regarding synthetic graphs, square grids seem challenging to solve because they have a relatively large burning number. Over these instances, although Gr could not find optimal solutions, GrP could. Nevertheless, the maximum execution time for Gr was 1587 seconds ($\sim$26.5 minutes) over a grid of order 102400, while GrP required 32606 seconds ($\sim$9.1 hours) over a grid of order 10000. Gurobi became unpractical on grids of order at least 1600 (See Table 4). By using Gurobi, Algorithm 4 corroborated all the previously known optimal solutions and found six new ones (tvshow, government, crocodile, DD6, DD349, and DD687); the implemented ILP (Expressions (8) to (11)) showed a better performance than the state-of-the-art mathematical formulations for GBP (See Tables 2 to 4).
@@ 7 Conclusions
This paper shows how GBP can be reduced to a series of CMCPs (See Proposition 2, Algorithm 1, and Expressions (8) to (11)). Through experimental results, we show that a deterministic greedy $(1/2)$-approximation algorithm for CMCP (Algorithm 2) can be used to find feasible solutions for GBP. Although the deterministic greedy heuristic for GBP, Gr (Algorithm 3), might not find optimal solutions of paths and cycles, it might have interesting theoretical properties over other graph families; we leave that research for future work. Nevertheless, the proposed heuristic performed unexpectedly well over real-world and synthetic graphs, finding most of the optimal and best-known solutions in reasonable amounts of time.

### v3

@@ 6 Discussion
GBP is an NP-hard minimization problem with application to information spreading problems. Some heuristics and metaheuristics have been proposed to find feasible solutions. However, research on more straightforward greedy approaches wasn’t fully addressed. This paper introduces a simple deterministic greedy heuristic, Gr (Algorithm 3), tailored for the CMCP, a problem to which GBP is reduced if $b(G)$ is known in advance; the issue of not knowing this value is tackled by using binary search (Algorithm 4). Although Gr cannot guarantee to find optimal solutions for paths and cycles for which exact $\mathcal{O}(n)$ algorithms exist, it performed unexpectedly well over real-world and synthetic graphs. Namely, for a benchmark dataset of real-world networks with order between 34 and 5908, Gr found 37 optimal solutions out of 47. Its plus version, GrP, found 44 out of 47 optimal solutions (See Table 2). Regarding larger real-world networks, Gr found 15 out of 15 optimal solutions; for most of these graph instances, GrP was very inefficient (See Table 3). Regarding synthetic graphs, square grids seem challenging to solve because they have a relatively large burning number. For these instances, although Gr could not find any optimal solution, GrP could. Nevertheless, the maximum execution time for Gr was 3282 seconds ($\sim$55 minutes) over a grid of order 102400, while GrP required 22301 seconds ($\sim$6.2 hours) over a grid of order 10000. Gurobi became unpractical on grids of order at least 4900 (See Table 4).
Using Gurobi for solving ILP-COV, all the previously known optimal solutions were corroborated, and six new ones were found (tvshow, government, crocodile, DD6, DD349, and DD687); ILP-COV showed a better performance than the mathematical formulations for GBP reported by [17] (See Tables 2 to 4).
@@ 7 Conclusions
This paper shows how GBP can be reduced to a series of CMCPs (See Proposition 2, Algorithm 1, and Expressions (7) to (10)). Through experimental results, we show that a deterministic greedy $(1/2)$-approximation algorithm for CMCP (Algorithm 2) can be used to find feasible solutions for GBP. Although the deterministic greedy heuristic for GBP, Gr (Algorithm 3), might not find optimal solutions of paths and cycles, it might have interesting theoretical properties over other graph families; we leave that research for future work. Nevertheless, the proposed heuristic performed unexpectedly well over real-world and synthetic graphs, finding most of the optimal and best-known solutions in reasonable amounts of time. Finally, the relationship between the GBP and the CMCP might be further exploited in the future, leading to more efficient algorithms.